OpenStudy (anonymous):

Kelvin Briley began a 114-mile bicycle trip to build up stamina for a triathlete competition. Unfortunately, his bicycle chain broke, so he finished the trip walking. The whole trip took 6 hours. If Kelvin walks at a rate of 4 miles per hour and rides at 24 miles per hour, find the amount of time he spent on the bicycle. Someone please show me how to do these type of word problems.

OpenStudy (anonymous):

whoever helps me, please don't give me the answer. Just help me step by step until I get the answer

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (dan815):

let x be time walking let y be time cycling v=d/t, so d=v*t 4x+24y=114 x+y=6

OpenStudy (dan815):

2 eqn, 2 unknown solve